In 2016, Teller Company sold 3,000 units at $600 each. Variable expenses were $420 per unit, and fixed expenses were $270,000. The same selling price, variable expenses, and fixed expenses are expected for 2017. What is Teller's break-even point in sales dollars for 2017?

A) $900,000

B) $2,700,000

C) $1,800,000

D) $2,571,429
Answer #1

Contribution margin=Sales-Variable cost  

=(600-420)=$180 per unit

Contribution margin ratio=Contribution margin/Sales

=(180/600)=0.3

Breakeven=Fixed expenses/Contribution margin ratio

=270,000/0.3

=$900,000